OpenStudy (melissa_something):

Find the unknown length in the right triangle. Simplify the answer if necessary. <3

OpenStudy (melissa_something):



OpenStudy (melissa_something):

\[a) \sqrt{10}\]

OpenStudy (melissa_something):

\[b) 2\sqrt{10}\]

OpenStudy (melissa_something):

c) 2

OpenStudy (melissa_something):

d) 58

OpenStudy (campbell_st):

well using pythagoras' theorem you have \[b^2 = 7^2 - 3^2\] so that's the start, the answer is a raducal, so you will need to simplify it

OpenStudy (anonymous):

is it b or 6?

OpenStudy (melissa_something):

It's b. Thanks, but I don't get if it will be B^2 =4^2?

OpenStudy (melissa_something):

like how does it even turn into a radical lol

OpenStudy (anonymous):

\[b2=7^2-3^2=49-9=40=4\times10,b=?\]